A genetic experiment involving peas yielded one sample of offspring consisting of 438 green peas and 173 yellow peas. Use a 0.05 significance level to test the claim that under the same circumstances, 26% of offspring peas will be yellow. Identify the null hypothesis, alternative hypothesis, test statistic, P-value, conclusion about the null hypothesis, and final conclusion that addresses the original claim.

Use the P-value method and the normal distribution as an approximation to the binomial distribution.

What are the null and alternative hypothesis?

What is the test statistic?

What is the P-value? What is the conclusion about the null hypothesis?

What is the final conclusion?
Answer #1

Answer:

Given,

sample proportion p^ = x/n = 173/438 = 0.395

Ho : p = 0.26

Ha : p != 0.26

test statistic z = (p^ - p)/sqrt(pq/n)

substitute values

= (0.395 - 0.26)/sqrt(0.26(1-0.26)/438)

= 6.44

P value = 0

Here we observe that, p value < alpha, so we reject Ho.

So there is sufficient evidence to support the claim.